OpenStudy (anonymous):

Write the equation in standard from (-4,3) m=1/2

OpenStudy (midhun.madhu1987):

y - y1 = m(x - x1) y - 3 = 1/2(x + 4) Simplify and express this equation in the Standard Form which is given by: Ax + By = C

OpenStudy (anonymous):

@midhun.madhu1987 I got it in slope intercept form y=1/2x+5 so how would I get it in standard form???

OpenStudy (midhun.madhu1987):

 Hope now u can convert it to the standard form...